A competitive exam was held in which five students took part namely Billy, Gerry, Clark, Peeta and Jonathan. In this exam, they had to answer five questions each out of which, two had multiple choices as a, b or c and three were simple true and false questions. The five of them gave different answers to the questions and the details are as given below.

Name 1 2 3 4 5

Gerry c b True True False

Billy c c True True True

Clark a c False True True

Peeta b a True True False

Jonathan a b True False True

None of the two students gave the same number of correct answers. Can you find out the correct answers to the question? Also, calculate the individual score of all the five guys.

Grade:12

2 Answers

Navjyot Kalra
10 years ago
It has been clearly stated that none of the two students gave same number of correct answers which leads us to two different possibilities of correct answers: Either (1 + 2 +3 + 4 + 5 = 15) or (0 + 1 + 2 + 3 + 4 = 10) stands true. After analyzing the question, the maximum number of correct answers possible with the answers Gerry, Billy, Clark, Peeta and Jonathan gave: Question 1 = 2 (b or c) Question 2 = 2 (b or c) Question 3= 4 (True) Question 4 = 4 (True) Question 5 = 3 (True) Let us first assume with the maximum number of correct answers possible according to our data which are 15 (2+2+4+4+3). According to it, Billy must have given all correct answers as he was the only one who answered True for questions 3, 4 and 5. But in that case, Gerry and Clark must have given exactly three correct answers. Furthermore, Peeta and Jonathan must have given two correct answers. Therefore none of them got all the five correct answers which is wrong according to what we have analyzed. Thus we move on to our next assumption according to which, the total number of correct answers are 10 (0+1+2+3+4). Now we must acknowledge that Questions 3 as well as Question 4 cannot be False as it will mean that the number of correct answers would not be 10. So the students giving wrong answers cannot be Gerry, Billy and Peeta. Assume that Jonathan got all wrong, it will suggest that Gerry, Clark and Peeta each would have at least two correct answers. Then, Billy would have to be the person with only one correct answer and in that situation, the correct answers for questions 1 and 2 would be a and a respectively. Since that is not possible as then the total number of correct answers would be 1 (a) + 1 (a) + 1 (False) + 4 (True) + 2 (Flase) = 9. Thus it must be Clark who has given all wrong answers. The correct answers are then b, a, True, False and False. Also, the scores are Clark (0), Billy (1), Gerry (2), Jonathan (3) and Peeta (4).
sai ramana
10 years ago
How does Jonathan`s correct number of answers become 3 from your correct answers list ( baTFF ). It should be 2 and contradicting the only effective condition and clue of the problem.